# Multiple Modes of Motion for the Effectiveness of Outer Hair Cells at High Frequencies

### v2

@@ Introduction
The mammalian hearing range may extends up to 100 kHz or beyond [1], depending on the species, quite remarkable for a biological system. Such function may call for a special rapid mechanism. Indeed, it is shown that the motility of outer hair cells (OHCs) based on “electromotility,” a kind of piezoelectricity, with which their cylindrical cell bodies are driven by the receptor potential, is essential for the sensitivity and frequency selectivity of the mammalian ear [2]. In addition, how OHCs are integrated into the rest of the cochlea is a critical issue, which requires extensive study to understand the mechanical basis of the mammalian hearing [3, 4].
@@
Some of the earlier theories treated the activity of OHCs as an additional pressure applied to the basilar membrane [5] or impedance of the basilar membrane [6, 7]. Some other treatments placed OHCs between two inertial masses, resulting in multiple degrees of freedom [8, 9, 10, 11]. The goal of these theories has been to explain the tuning curve of the ear.
@@
Now, let us focus on the issue of the upper bound of auditory frequency. One of the difficulties of explaining the effectiveness of OHCs was the low-pass nature of their intrinsic electric circuit. This problem can be addressed by introducing piezoelectric resonance [12, 13]. However, the model employed in the treatment assumed that soft OHCs and the stiff basilar membrane (BM) are components of the same oscillator. The upper limit obtained was $\sim$10 kHz, much higher than the roll-off frequency (up to several kHz) of the intrinsic electric circuit, but still not high enough for covering the auditory frequency [13].
@@
A possible solution to this problem is that the OHCs and the basilar membrane are elements of separate oscillators, and those oscillators are coupled so that energy can be transferred between them. Multiple modes of motion appears to be consistent with recent experimental reports, including those with optical coherence tomography (OCT) [14]. Those reports show complex motion of the inner ear, which depends on the stimulation intensity as well as frequency [15, 16, 17, 18].
@@
Here we examine a set of two oscillators, heavy and light, coupled with either elastic or viscous element as the simplest examples of a system with multiple modes of motion. As a control, a single mode oscillator with an OHC is described to illustrate the effect of impedance mismatch. Then systems of two coupled oscillators are examined. The light oscillator (LO) incorporates an OHC and the heavy one (HO) includes the basilar membrane. Coupling is either elastic or viscous. The OHC is stimulated by either the motion of HO or that of LO.
@@
The present treatment focuses on local energy balance, and intended to address the upper bound of the auditory frequency. Such a location may correspond to the basal end of the cochlea, where traveling waves start.
@@
Let us start from a simple model oscillator system, in which an OHC incorporated as illustrated in Fig. 1. Here we follow the one-dimensional model [13], rather than the more realistic cylindrical model [19] for simplicity.
First, we drive the system by an external force $F$. The equation of motion can be written down as
$\left(m\frac{d^{2}}{dt^{2}}+\eta\frac{d}{dt}+k_{e}\right)X=k_{o}(X_{\infty}-X)+F,$ (1)
where $m$ is the mass, $\eta$ drag coefficient, $k_{e}$ the stiffness of the external elastic load, $k_{o}$ the material stiffness of the OHC, $X$ is the length of the OHC, and $k_{e}$ the stiffness of external load. $X_{\infty}$ is the length of the cell, which satisfies the equilibrium condition for the voltage $V$ at the given moment.
Figure 1: A single oscillator with an OHC. The oscillator consists of mass $m$, an external elastic element with stiffness $k_{e}$, and a damper with drag coefficient $\eta$ and an outer hair cell (OHC), which responds to the movement of mass (dashed red arrow). This system is driven by a sinusoidal waveform with angular frequency $\omega$ (wavy blue arrow).
If the force applied is a periodic waveform with angular frequency $\omega$, we can write
@@
Then we have
$x\left(-\overline{\omega}^{2}+i\frac{\overline{\omega}}{\overline{\omega}_{\eta}}+1+\frac{ku_{a}}{k_{e}+k_{o}}\right)=-\frac{k_{o}}{k_{e}+k_{o}}u_{b}v+\frac{f}{k_{e}+k_{o}},$ (3)
where $\overline{\omega}(=\omega/\omega_{r})$ is the frequency normalized to the mechanical resonance frequency $\omega_{r}=\sqrt{(k_{e}+k_{o})/m}$, and $\overline{\omega}_{\eta}$ is normalized viscoelastic roll-off frequency. Parameters $u_{a}$ and $u_{b}$ depend on the piezoelectric properties of the OHC (See Table 1). For derivation, see Appendix A.
Now, consider that the voltage waveform is generated by changes of hair bundle resistance $\hat{r}$ due to bending. The effect of hair bundle resistance $R_{a}$ on the membrane potential $V$ can be expressed
$(e_{ec}-V)/R_{a}=(V-e_{K})/R_{m}+C_{0}\;dV/dt-Nq\;dP/dt,$ (4)
where $e_{ec}$ is the endocochlear potential, $e_{K}$ is the resting potential of OHC, which is primarily determined by K+ conductance, $R_{a}$ hair bundle conductance, $C_{0}$ the voltage-independent membrane capacitance of the cell, $q$ and $N$, respectively, the unit charge the motile element transfers across the membrane and the number of such motile elements in the cell. $P$ is the fraction of such motile elements in the extended state, which contributes to elongation of the cylindrical cell.
With periodic stimulation, the hair bundle conductance $R_{a}$ consists of a time independent component $R_{0}$ and a sinusoidal component with relative amplitude $\hat{r}$. Thus the receptor current with frequency $\omega$ is expressed by
$-i_{0}\hat{r}=(\sigma+i\omega C_{0})v-i\omega Nqp.$ (5)
Here $i_{0}=(e_{ec}-e_{K})/(R_{0}+R_{m})$ is the steady state current and $\sigma=1/R_{0}+1/R_{m}$ the steady state conductance, which can be dropped for high frequency stimulation because it is smaller than $\omega C_{0}$. The quantity $p$ is the amplitude of $P$ with frequency $\omega$.
Thus the variable $v$ can be replaced by $\hat{r}$,
@@
$\hat{f}$ $=f/(k_{e}+k_{o}).$ (7c)
Here $u$, $u_{a}$, and $u_{q}$ are quantities that represent hair bundle sensitivity and piezoelectricity. See Table 1 for definitions (for derivations see Appendix A). The quantity $\gamma$ represents relative voltage sensitivity of prestin, the motile membrane protein of OHCs [20]. The maximal value is set to unity (i.e. $0\leq\gamma\leq 1$). It will be referred to as the “operating point parameter. ” Notice also that both $A$ and $B$ depend on the ratio $k_{e}/k_{o}$ of external load to the intrinsic stiffness of the OHC.
The first term of the right-hand side of Eq. 6, which contains $A$, works as an amplifying term, counteracting the drag term on the left-hand side. However, it is inversely proportional to the frequency $\omega$, unlike negative drag. The reason for this frequency dependence stems from the receptor potential (Eq. 5), where capacitive current overwhelms resistive current at high frequencies. The parameter $B$ represents the contribution of the motile element to the stiffness of the cell body of the OHC.
Table 1: Parameter definitions. $k_{B}$ Boltzmann’s constant, $T$ the temperature. The motile function of OHCs is described by a two state model. $P_{0}$ is the fraction of one of the two states in a single cell at the operating point. ⋆⋆ The definition of $\kappa$ given here assumes the presence of $k_{c}$, the elastic element between the two oscillators. In the absence of such coupling, $k_{c}=0$. See Appendix A for derivations.
@@
$P_{0}$ motor conformation $\gamma$ $4P_{0}(1-P_{0})$
$\beta$ $1/(k_{B}T)$ $u_{a}$ $\beta a^{2}k_{o}N/4$
$a$ prestin displacement $u_{b}$ $\beta aqN/4$
$q$ prestin charge $u$ $\beta i_{0}aqN/(4C_{0})$
$N$ number of prestin $u_{q}$ $\beta q^{2}N/(4C_{0})$
$C_{0}$ cell capacitance $\kappa$⋆ $k_{o}/(k_{o}+k_{e}+k_{c})$
$i_{0}$ resting current $\omega_{r}$ resonance frequency
@@
Even though we have assumed that the amplitude $x$ is small, Eq.  8 may leads to a large $x$, depending on the parameter values. In such cases, such an amplitude should be an artifact of linearization. The main reason is that the coefficient $g$ must be smaller for a large $x$ due to nonlinearity of the response function. In addition, the factors $A$ and $B$ are also affected because they depend on $\gamma$, which is determined by the operating point of electromotility.
@@
Consider a system with two harmonic oscillators, light and heavy. The light oscillator (LO) consists of an OHC and an elastic load $k_{e}$ and inertia $m$. The displacement of this oscillator is $X$. The heavy oscillator (HO) consists of the basilar membrane and inertia $M$, an elastic element with stiffness $K$, and viscous load with drag coefficient $\eta$. We can assume in general that the two oscillators are coupled by an elastic element with stiffness $k_{c}$ and a viscous element with drag coefficient $\eta_{c}$, and external force $F$ is applied to the HO. The motion of the two oscillators are described by,
$\left(m\frac{d^{2}}{dt^{2}}+\eta_{1}\frac{d}{dt}+k_{e}+k_{o}+\gamma B\right)X$ $=i\frac{\gamma A}{\overline{\omega}}\;\hat{r}+\left(k_{c}+\eta_{c}\frac{d}{dt}\right)(Y-X)$ (10a)
@@
where quantities $A$ and $B$ are defined by Eq. 7 with viscous coupling but elastic load modifies these quantifies. That means
$A$ $=(u/\omega_{r})\;k_{0}/(k_{e}+k_{o}+k_{c})$ (11a)
$B$ $=u_{a}(k_{e}+k_{c})/(k_{e}+k_{c}+k_{o})+u_{q}.$ (11b)
For this reason, the quantities $A$ and $B$ depends on $k_{c}$ even in the case where $k_{e}$ is kept constant.
An issue for this model is which oscillator include the presence of the main drag with coefficient $\eta$. Here we assume that the main drag belongs to HO because the HO is heavier than the LO. In the case we assume that the main drag belong to LO, it turns out that the OHC only amplifies the amplitude of LO and it cannot affect the oscillation of HO.
OHC works on the LO, its hair bundle sensing the displacement $X$ of LO or the displacement $Y$ of the HO, or both. For simplicity, let us consider only two cases, introducing the hair bundle sensitivity $g$ to displacements: $\hat{r}=gX$, where the hair bundle senses the displacement of LO, and $\hat{r}=gY$ where the hair bundle senses the displacement of HO.
Classical analyses show that energy transfer between two coupled oscillators is rather complex despite the apparent simplicity of the equations [21, 22]. For simplicity, here we only examine only the response to continuous sinusoidal stimulation and separate each displacement into a constant component and time dependent component.
@@
where $\omega$ is the angular frequency of stimulation.
@@
In general, bending of the OHC hair bundle may depend on $x$, the displacement of the LO as well as $y$, the displacement of the HO. The simplest assumption to represent such dependence would be to express $\hat{r}$ as a linear combination of $x$ and $y$. For simplicity, two extreme cases will be examined at first: the case where $\hat{r}$ depends only on $y$, and another where it depends only on $x$. Then we proceed to a more general case, where $\hat{r}$ depends on the combination of $x$ and $y$.
Since we have so many parameters, we reduce the number of parameters by considering a number of separate cases. First, coupling includes viscous and elastic components. However, we assume coupling is purely elastic or purely viscous. Second, we further assume that the hair bundle of the OHC is stimulated either by the displacement of the HO, which incorporates the basilar membrane, or the displacement of the LO, which includes the OHC. For brevity, let us call the former as “HO-driven mode” and the latter “LO-driven mode.”
If we assume that subtectorial drag is the dominant one, the drag coefficient of which is $\eta$, the HO-driven case includes the subtectorial drag because both depend on displacement $y$. Thus, $\eta_{2}=\eta$ and $\eta_{1}=0$. In LO-driven mode, likewise $\eta_{2}=0$ and $\eta_{1}=\eta$ because they are both associated with $x$, the displacement of the LO.
@@
We assume that the main drag of the system belongs to HO, and that the dominant drag is due to the shearing motion in the subtectorial space, then a natural assumption would be that the hair bundle of the OHC is stimulated by the motion of the HO. Then, we may express changes of the hair bundle conductance by $\hat{r}=gY$ with the sensitivity $g$.
@@
Let us start from the system with viscous coupling with characteristic frequency $\omega_{\eta}$ on the heavy oscillator. Given the stiffness ratio $s$ of the heavy oscillator to the light oscillator, the characteristic frequency on the light oscillator is $\omega_{\eta}/s$. Introducing new parameters defined by $f=f_{B}/K$ and $s=\!K/k$, the set of equations can be written as
@@
Elastic coupling introduces an elastic element $K_{c}$ in each of the equation of motion, elevating the resonance frequency more significantly in LO than in HO. Since energy transfer between the oscillators depends on the difference in the resonance frequencies of the two oscillators, adjustment should be made to minimize the difference in the resonance frequency.
@@
introducing new parameters $c$ and $\overline{\omega}_{1}$. Of these parameters, $c$ is defined by $c=\!k_{c}/K$. We may assume $c<1$ because the coupling should not be too strong.
The parameter $\overline{\omega}_{1}$ is to eliminate the dependence of the resonance frequency of LO by the elastic coupling element, which elevates the resonance frequency. This effect is less significant for the HO because $c$ is small. However, it is quite pronounced for the LO if the ratio $s$ is large. This effect should be compensated because a larger difference in the resonance frequencies of the two oscillators decreases energy transfer between the oscillators. The parameter $\overline{\omega}_{1}$, in effect, allows adjustment of the mass of the LO so that energy transfer is optimized for a large values for the ratio $s$.
@@
An alternative assumption regarding OHC stimulation is that the displacement $x$ of the LO affects the bending of OHC hair bundle, creating a direct feedback loop.
@@
Eq. 17 indicates that the LO reduces, rather than increases, the amplitude of the HO if the effect of OHC is large. If the term $Ag/\overline{\omega}$ dominates over $cs$, Eq. 15a makes the phase of $x$ delayed from that of $y$ up to $-\pi/2$. In addition, the magnitude of $x$ is expected to be larger near the resonance frequency ($\overline{\omega}\approx 1$) because the coefficient of $x$ in Eq. 15a can be much smaller than the coefficient of $y$. This feature will be illustrated later with numerical examination.
General cases
OHCs can be stimulated by a combination of the displacement of LO and that of HO. In addition, the coupling of these two oscillators includes both viscous and elastic elements. Such cases are not examined here because those combinations introduce additional parameters, which are needed to describe relative significance of these terms.
@@
Given the number of parameters in our set of equations, our numerical analysis is to survey conditions, under which the OHC is more effective in coupled oscillators than in the single mode oscillator as the amplifier.
In the following, values for $A$ and $B$, which represent the activity of OHCs are examined using experimental data available in a best frequency range. Then the amplitude of the single mode model is evaluated. After that, the amplitudes of coupled oscillator models are evaluated. Finally, the effectiveness of the coupled oscillator models in higher frequency ranges will be discussed as an extrapolation.
For examining the performance of coupled oscillator models, experimental values obtained from guinea pigs near 20 kHz location are available. See Table 2.
@@
Since $A=(u/\omega_{r})\kappa$ and $B=u_{a}(1-\kappa)+u_{q}$ as defined earlier by Eq. 7, $B$ is approximately 2. It is not sensitive to the load ratio $\kappa$ because it $u_{a}$ is smaller than $u_{q}$.
@@
Let us assume that the characteristic stiffness of HO is that of the basilar membrane (BM). At the location of 20 kHz best frequency in guinea pigs is 3 mm from the stapes according to the Greenwood function, assuming the length of the BM is 18.5 mm [28, 29]. The stiffness of the location is about 0.21 N/m measured with a probe with 25 $\mu$m diameter tip [30]. This value is compatible with the ones obtained with a probe with 10 $\mu$m tip [31]. Thus, the stiffness ratio $s_{o}$ is $\sim$10.
@@
Of OHC parameters, $B$ does not depend heavily on the resonance frequency because it is dominated by the electrical term $u_{q}$ (See Eq. 7b). The amplification parameter $A$ decreases at higher frequency locations because it is inversely proportional to the resonance frequency $\omega_{r}$. However, such reduction is compensated by other factors. OHCs at higher frequency locations have larger resting current $i_{0}$ owing to larger hair bundle conductance and larger structural stiffness $k_{o}$ of the cell body owing to its shorter cell length (See Eq. 7a and Table 1).
@@
Figure 5: Amplitude of single mode oscillator plotted against frequency (normalized to $\omega_{r}$). Stiffness ratio of load to OHC is 1, 1.2, and 5. $\gamma$ (operating point variable) runs from 0.03 (left) to 0.24 with 0.03 increment. The unit of the ordinate axis is $f(=F/K)$.
Let us start from the performance of the single mode oscillator before examining coupled oscillator models. The steady state amplitude is obtained by solving Eq. 8. The amplitude is normalized by the input $\hat{f}$ (Figs. 5. As $\gamma$, the operating point variable of the OHC, increases, the peak amplitude shifts to higher frequencies independent of the elastic load (Fig. 6B)). The peak height sharply increases with $\gamma$ if the elastic load is relatively small, i.e. the ratio $k_{e}/k_{o}$ is 1 or smaller (Fig. 6A). However, the amplitude does not increase with $\gamma$ even if $k_{e}/k_{o}$ is as small as 1.2. With larger elastic load, such as $k_{e}/k_{o}\geq 5$, the peak amplitude does not increase at all.
@@
In the following, conditions for large gain in amplitude are sought in all four cases. Those cases are, HO driven modes with elastic coupling and viscous coupling, and LO driven modes with elastic and viscous coupling. For each condition, a single example is given. These examples are typical ones but they do not necessarily show all shared features of the conditions.
@@
It would be natural to start from HO-driven mode because we assumed that HO is associated with the dominant drag, which likely stems from the shear in the subtectorial space and this shear should stimulate the hair bundle of the OHC.
@@
Here, we assume that the coupling of the two oscillators is elastic. It is possible to find conditions, under which the amplitude of the LO reaches $\sim$20 fold larger than the amplitude of the single mode oscillator (Fig. 8A). Here the abscissa is the operating point variable, which depends the voltage. It reaches unity at the optimal operating voltage. The amplitude of HO is about a half of that of LO. The peak frequency is about the same for the two oscillators and increases with the operating point variable (Fig. 8B).
@@
Figure 8: HO-driven elasticity coupled oscillators (HOE). A: Amplitude gain over SO mode, B: Peak frequency, and C: The phase with respect to the external force, of each oscillator are plotted against $\gamma$. LO (red) and HO (blue). D: The ratio of power output to power input. The set of parameter values: $c$=0.2, $\omega_{1}$=0.29 in Eq. 15.
@@
Here we assume that the OHC is primarily driven by the motion of LO rather than that of HO to examine if the amplitude gain can be larger. The meaning of this mode will be discussed later. It turns out that LO-driven modes are more effective than HO-driven modes.
@@
The amplitude gain for both LO and HO increases with the operating point variable $\gamma$ (Fig. 9A). HO has somewhat higher peak frequency than LO for small $\gamma$ (Fig. 9B). The phases of HO and O are both close to $\pi/4$ and HO tends to be ahead of LO except for where $\gamma$ is small (Fig. 9C). The power amplification is not very large, similar to the amplitude of HO.

### v3

@@ Abstract
Significance
The mammalian ear depends on outer hair cells as the cochlear amplifier. To perform this biological function, soft outer hair cells need to amplifying the vibration of the much stiffer basilar membrane, overcoming the stiffness mismatch. The present paper shows that a system of two coupled oscillator, the simplest of more complex systems, allows effective transmission of the power generated by outer hair cells to the vibration of the basilar membrane. Therefore, the elaborate structure of the inner ear, which supports multiple modes of motion, must be a key to the exquisite performance of the mammalian ear. In addition, the properties of outer hair cells obtained from isolated cells preparations are compatible with their physiological function.
@@ Introduction
The mammalian hearing range may extend up to 100 kHz or beyond [1], depending on the species, quite remarkable for a biological system. Such function may call for a special rapid mechanism. Indeed, it is shown that “electromotility” of outer hair cells (OHCs), a kind of piezoelectricity with which their cylindrical cell bodies are driven by the receptor potential, is essential for the sensitivity and frequency selectivity of the mammalian ear [2]. In addition, how OHCs are integrated into the rest of the cochlea is a critical issue, which requires extensive study to understand the mechanical basis of the mammalian hearing [3, 4].
@@
Some of the earlier theories treated the activity of OHCs as an additional pressure applied to the basilar membrane (BM) [5] or impedance of the BM [6, 7]. Some other treatments placed OHCs between two inertial masses, resulting in multiple degrees of freedom [8, 9, 10, 11, 12]. The goal of these theories has been to explain the tuning curve of the ear.
@@
Now, let us focus on the issue of the upper bound of auditory frequency. One of the difficulties of explaining the effectiveness of OHCs was the low-pass nature of their intrinsic electric circuit. This problem can be addressed by introducing piezoelectric resonance [13, 14]. However, the model employed in the treatment assumed that soft OHCs and the stiff BM are components of the same oscillator. The upper limit obtained was $\sim$10 kHz, much higher than the roll-off frequency (up to several kHz) of the intrinsic electric circuit, but still not high enough for covering the auditory frequency [14].
@@
A possible solution to the limitation imposed by impedance mismatch is that the OHCs and the BM are elements of separate oscillators, and those oscillators are coupled so that energy can be transferred between them. Multiple modes of motion appears to be consistent with recent experimental reports, including those with optical coherence tomography (OCT) [15]. Those reports show complex motion of the inner ear, which depends on the stimulation intensity as well as frequency [16, 17, 18, 19].
Although recent studies have shown that the neuronal output of the cochlea is more closely correlated to the motion of the reticular lamina rather than that of the BM [20, 21], BM motion is essential because it is the basis for the tonotopic map of the cochlea [22]. The objective of the present study is to examine the performance of simple cases of coupled oscillators, which retain the essential features of the cochlea, using the properties of OHCs determined from isolated cell preparations.
@@
We examine a set of two oscillators, heavy and light, coupled with either elastic or viscous element as the simplest examples of a system with multiple modes of motion. At first, a single mode oscillator with an OHC is described to illustrate the effect of impedance mismatch. Then systems of two coupled oscillators are examined. The light oscillator (LO) incorporates an OHC and the heavy one (HO) includes the BM. Coupling is either elastic or viscous. The OHC is stimulated by either the motion of HO or that of LO.
@@
The present treatment focuses on local energy balance, and intended to address the upper bound of the auditory frequency, even though the condition of local energy balance applies to any location, where traveling waves of a given frequency stops [23].
@@
Let us start from two simple model oscillator systems, in which an OHC is incorporated (Fig. 1). The equation of motion can be formally written as
$\left(m\frac{d^{2}}{dt^{2}}+\eta\frac{d}{dt}+k_{o}+k_{e}\right)X=F_{\mathrm{OHC}}+F,$ (1)
where $m$ is the mass, $\eta$ drag coefficient, $k_{e}$ the stiffness of the external elastic load, $k_{o}$ the material stiffness of the OHC, $X$ is the length of the OHC, and $k_{e}$ the stiffness of external load. $F$ is an external force and $F_{\mathrm{OHC}}$ is the force generated by the OHC.
OHC force $F_{\mathrm{OHC}}$ is generated by the difference of mechanoelectric coupling free energy between the given moment and the values of for model illustrated by Fig. 1 can be expressed by the free energy difference [24, 14, 25].
Figure 1: A single oscillator model. The oscillator consists of mass $m$, an external elastic element with stiffness $k_{e}$, and a damper with drag coefficient $\eta$ and an outer hair cell (OHC), which responds to the movement of mass (dashed red arrow). This system is driven by a sinusoidal waveform with angular frequency $\omega$ (wavy blue arrow).
Consider the case, where a force $F$ applied has a periodic waveform with angular frequency $\omega$, we can write
@@
The derivation of the equation for the single oscillator model follows previous papers [24, 14, 25] as shown in Appendix. The basic assumption underlying the equation is that the driving force of the OHC at a given moment is due to the free energy difference from the equilibrium state for the given voltage and the force applied to the motile element of the OHC.
The equation of motion can be expressed using changes in hair bundle resistance $\hat{r}$ instead of voltage changes $v$, as follows. See Appendix for details of the derivation.
@@
$\hat{f}$ $=f/(k_{e}+k_{o}),$ (4c)
and $\overline{\omega}(=\omega/\omega_{r})$ is the frequency normalized to the mechanical resonance frequency $\omega_{r}=\sqrt{(k_{e}+k_{o})/m}$, and $\overline{\omega}_{\eta}$ is normalized viscoelastic roll-off frequency.
Quantities $u$, $u_{a}$, and $u_{q}$ represent hair bundle sensitivity and piezoelectricity of the OHC. More specifically, $u$ represents mechanoelectric coupling, $u_{a}$ mechanical term, and $u_{a}$ electrical term, while all of them depend on hair bundle sensitivity. See Table 1 for definitions (for derivations see Appendix). The quantity $\gamma$ represents relative voltage sensitivity of prestin, the motile membrane protein of OHCs [26]. The maximal value is set to unity (i.e. $0\leq\gamma\leq 1$). It will be referred to as the “operating point parameter.” Notice also that both $A$ and $B$ depend on the ratio $k_{e}/k_{o}$ of external load to the intrinsic stiffness of the OHC.
The first term of the right-hand side of Eq. 3, which contains $A$, works as an amplifying term, counteracting the drag term on the left-hand side. However, it is inversely proportional to the frequency $\omega$, unlike negative drag. The reason for this frequency dependence stems from the receptor potential, where capacitive current overwhelms resistive current at high frequencies. The parameter $B$ represents the contribution of the motile element to the stiffness of the cell body of the OHC.
Table 1: Parameter definitions
@@
$P_{0}$ motor conformation $\beta$ $1/(k_{B}T)$
$a$ prestin displacement $\gamma$ $4P_{0}(1-P_{0})$
$\omega_{r}$ resonance frequency $u_{a}$ $\beta a^{2}k_{o}N/4$
$q$ prestin charge $u_{b}$ $\beta aqN/4$
$N$ number of prestin $u$ $\beta i_{0}aqN/(4C_{0})$
$C_{0}$ cell capacitance $u_{q}$ $\beta q^{2}N/(4C_{0})$
$i_{0}$ resting current $\kappa$⋆ $k_{o}/(k_{o}+k_{e}+k_{c})$
@@
$k_{c}$ coupling spring $f_{B}$ external force (amplitude)
$\omega_{1}$ resonance frequency ratio $\omega_{\eta}$ $K/\eta$
$\hat{r}$ relative change of hair bundle resistance
Note: $k_{B}$ is Boltzmann’s constant, $T$ the temperature. The motile function of OHCs is described by a two state model. $P_{0}$ is the fraction of one of the two states in a single cell at the operating point. ⋆⋆ The definition of $\kappa$ given here assumes the presence of $k_{c}$, the elastic element between the two oscillators. In the absence of such coupling, $k_{c}=0$. See Appendix for derivations.
@@
Consider a system with two harmonic oscillators, light and heavy. The light oscillator (LO) consists of an OHC and an elastic load $k_{e}$ and inertia $m$. The displacement of this oscillator is $X$. The heavy oscillator (HO) consists of the BM and inertia $M$, an elastic element with stiffness $K$, and viscous load with drag coefficient $\eta$. We can assume in general that the two oscillators are coupled by an elastic element with stiffness $k_{c}$ and a viscous element with drag coefficient $\eta_{c}$, and external force $F$ is applied to the HO (Figs. 2 and 3). The motion of the two oscillators are described by,
$\left(m\frac{d^{2}}{dt^{2}}+\eta_{1}\frac{d}{dt}+k_{e}+k_{o}\right)X$ $=F_{\mathrm{OHC}}+\left(k_{c}+\eta_{c}\frac{d}{dt}\right)(Y-X)$ (7a)
@@
where $F_{\mathrm{OHC}}$ is active force exerted by the OHC, being stimulated at its hair bundle.
An issue for this model is which oscillator includes the main drag with coefficient $\eta$. Here we assume that the main drag belongs to HO. If we assume that the main drag belong to LO, the OHC only amplifies the amplitude of LO and it cannot affect the oscillation of HO.
Classical analyses show that energy transfer between two coupled oscillators is rather complex despite the apparent simplicity of the equations [27, 28]. Here we only examine only the response to continuous sinusoidal stimulation.
@@
where $\omega$ is the angular frequency of stimulation. Notice here that notation $f_{B}$ is used for the amplitude of external force in the systems of coupled oscillators. Now $F_{\mathrm{OHC}}$ can be expressed in a manner similar to in Eq. 5 using $A$ and $B$. In the case of viscosity coupling these quantities are defined by Eq. 4. With elastic load coupling these quantifies are expressed by
$A$ $=(u/\omega_{r})\;k_{0}/(k_{e}+k_{o}+k_{c})$ (9a)
$B$ $=u_{a}(k_{e}+k_{c})/(k_{e}+k_{c}+k_{o})+u_{q}.$ (9b)
In general, bending of the OHC hair bundle that results in $\hat{r}$ may be proportional to $x$, the displacement of the LO as well as $y$, the displacement of the HO. For simplicity, two extreme cases will be examined here: $\hat{r}$ depends only on $y$ (HO-driven mode), or only on $x$ (LO-driven mode). Coupling may have both viscous and elastic components. However, we assume coupling is purely elastic or purely viscous for simplicity.
@@
The assumptions that the main drag of the system belongs to HO, and that the dominant drag is due to the shearing motion in the subtectorial space, may lead to hair bundle sensitivity to HO motion (Fig. 2), i.e. $\hat{r}=gY$.
@@
HO has viscoelastic roll-off frequency $\omega_{\eta}$. With new parameters defined by $s=\!K/k$ and re-defined external force amplitude $f=f_{B}/K$, the set of equations can be written as
@@
Coupling introduced by an elastic element $k_{c}$ would elevate the resonance frequency of LO more than that of HO. Since energy transfer between the oscillators depends on the difference in the resonance frequencies of the two oscillators, an parameter $\omega_{1}$ is introduced to adjust the resonance frequency of LO.
@@
where $c$ is defined by $c=\!k_{c}/K$.
@@
An alternative assumption regarding OHC stimulation is that the displacement $x$ of the LO affects the bending of OHC hair bundle, creating a direct feedback loop (Fig. 3).
@@
For numerical analysis of the performance of the coupled oscillators described, determining the values for parameters $A$ and $B$, OHC stiffness $k_{o}$ and BM stiffness $K$ is essential. For this purpose experimental values from guinea pigs near 20 kHz location are available. See Table 2.
@@
Since $A=(u/\omega_{r})\kappa$ and $B=u_{a}(1-\kappa)+u_{q}$ as defined earlier by Eq. 4, $B$ is approximately 2. It is not sensitive to the load ratio $\kappa$ because it $u_{a}$ is smaller than $u_{q}$ (Fig. 4).
@@
Assume that the stiffness $K$ of HO is from the BM. At the location of 20 kHz best frequency in guinea pigs is 3 mm from the stapes according to the Greenwood function, assuming the length of the BM is 18.5 mm [34, 35]. The stiffness of the location is about 0.21 N/m measured with a probe with 25 $\mu$m diameter tip [36]. This value is compatible with the ones obtained with a probe with 10 $\mu$m tip [37]. Thus, the stiffness ratio $s_{o}$ is $\sim$10.
@@
Of the OHC parameters, $B$ does not depend heavily on the resonance frequency (Fig. 4) because it is dominated by the electrical term $u_{q}$ (See Eq. 4b). The amplification parameter $A$ decreases at higher frequency locations because it is inversely proportional to the resonance frequency $\omega_{r}$. However, such reduction is compensated by other factors. OHCs at higher frequency locations have larger resting current $i_{0}$ owing to larger hair bundle conductance and larger structural stiffness $k_{o}$ of the cell body owing to its shorter cell length (See Eq. 4a and Table 1).
@@
Figure 5: Amplitude of single mode oscillator plotted against frequency (normalized to $\omega_{r}$). Peak values are re-plotted in Fig. 6A. Stiffness ratio of load to OHC is 1, 1.2, and 5. $\gamma$ (operating point variable) runs from 0.03 (left) to 0.24 with increment of 0.03. The unit of the ordinate axis is $f(=F/K)$.
Let us start from the performance of the single mode oscillator before examining coupled oscillator models. The steady state amplitude is obtained by solving Eq. 5. The amplitude is normalized by the input $\hat{f}$ (Figs. 5). As $\gamma$, the operating point variable of the OHC, increases, the peak amplitude shifts to higher frequencies independent of the elastic load (Fig. 6B). The peak height sharply increases with $\gamma$ if the elastic load is relatively small, i.e. the ratio $k_{e}/k_{o}$ is 1 or smaller (Fig. 6A). However, the amplitude does not increase with $\gamma$ even if $k_{e}/k_{o}$ is as small as 1.2. With larger elastic load, such as $k_{e}/k_{o}\geq 5$, the peak amplitude does not increase at all.
@@
Conditions for large gain in amplitude are sought in all four cases: HO driven modes with elastic coupling and viscous coupling, and LO driven modes with elastic and viscous coupling. For each condition, a single example is given. These examples are typical ones but they do not necessarily show all shared features of the conditions.
@@
It would be natural to start from HO-driven mode because we assumed that HO is associated with the dominant drag, which likely stems from the shear in the subtectorial space and this shear should stimulate the hair bundle of the OHC (Fig. 2).
@@
Here, we assume that the coupling of the two oscillators is elastic. It is possible to find conditions, under which the amplitude of the LO reaches $\sim$20 fold larger than the amplitude of the single mode oscillator (Fig. 8A). The abscissa is the operating point variable, which depends the voltage. It reaches unity at the optimal operating voltage. The peak frequency is about the same for the two oscillators and increases with the operating point variable (Fig. 8B).
@@
Figure 8: HO-driven elastically coupled oscillators (HOE). A: Amplitude gain over SO mode, B: Peak frequency, and C: The phase with respect to the external force, of each oscillator are plotted against $\gamma$. LO (red) and HO (blue). D: The ratio of power output to power input. The set of parameter values: $c$=0.2, $\omega_{1}$=0.29 in Eq. 12.
@@
Here we assume that the OHC is primarily driven by the motion of LO rather than that of HO (Fig. 3) to examine if the amplitude gain can be larger. The meaning of this mode will be discussed later. It turns out that LO-driven modes are more effective than HO-driven modes.
@@
The amplitude gain for both LO and HO increases with the operating point variable $\gamma$ (Fig. 9A). HO has somewhat higher peak frequency than LO for small $\gamma$ (Fig. 9B). The phases of HO and LO are both close to $\pi/4$ and HO tends to be ahead of LO except for where $\gamma$ is small (Fig. 9C). The power amplification is not very large, similar to the amplitude of HO.